please I need help solving this word problem.
one integer is 8 more than another. Their product is 19 more than their sum. what are the two integer?
Here is how is set up my fraction x(8+x)=19
8x+x^2=19
0=x^2+8x+19

Answer by mananth(16946)

x
x+8
product is 19 more than the their sum
sum is x+x+8
x(x+8) = 19 +(x+(x+8))
x^2+8x=19+2x+8
x^2+6x-27=0

x^2+9x-3x-27=0
x(x+9)-3(x+9)=0
(x+9)(x-3) = 0
x=-9 OR 3
-9,-1
OR
3,11